OpenStudy (anonymous):

8x - 1 = -4x + 23 x =

OpenStudy (anonymous):

First we will want to get x by itself on one side so... 8x - 1 = -4x + 23 add 4x 12x - 1 = 23 add 1 12x = 24 now divide by 12 and that is your answer :)

OpenStudy (anonymous):

oh my god this is simple thank you!

OpenStudy (imtant):

Bring the -4x to the right,while -1 to the left. 8x+4x=23+1 (the sign will become conversely) so you can solve it then. 12x=24 x=2..

OpenStudy (anonymous):

the answer would be 2

OpenStudy (anonymous):

thanks guys!!! :)))))

OpenStudy (anonymous):

Yep :)

OpenStudy (imtant):

Ok,wc..

OpenStudy (anonymous):

love u